# Learning Minimal NAP Specifications for Neural Network Verification

### v2

@@ Abstract.
Specifications play a crucial role in neural network verification. They define the precise input regions we aim to verify, typically represented as L-infinity norm balls. While recent research suggests using neural activation patterns (NAPs) as specifications for verifying unseen test set data, it focuses on computing the most refined NAPs, often limited to very small regions in the input space. In this paper, we study the following problem: Given a neural network, find a minimal (coarsest) NAP that is sufficient for formal verification of the network’s robustness. Finding the minimal NAP specification not only expands verifiable bounds but also provides insights into which neurons contribute to the model’s robustness. To address this problem, we propose several exact and approximate approaches. Our exact approaches leverage the verification tool to find minimal NAP specifications in either a deterministic or statistical manner. Whereas the approximate methods efficiently estimate minimal NAPs using adversarial examples and local gradients, without making calls to the verification tool. This allows us to inspect potential causal links between neurons and the robustness of state-of-the-art neural networks, a task for which existing verification frameworks fail to scale. Our experimental results suggest that minimal NAP specifications require much smaller fractions of neurons compared to the most refined NAP specifications, yet they can significantly expand the verifiable boundaries to several orders of magnitude larger.
Do, Not, Us, This, Code, Put, the, Correct, Terms, for, Your, Paper
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
@@ 1. Introduction
The growing prevalence of deep learning systems in decision-critical applications has elevated safety concerns regarding AI systems, such as their vulnerability to adversarial attacks  (Goodfellow et al., 2015; Dietterich and Horvitz, 2015). Therefore, the verification of AI systems has become increasingly important and attracted much attention from the research community. The field of neural network verification largely follows the paradigm of software verification – using formal methods to verify desirable properties of systems using rigorous mathematical specifications and proofs (Wing, 1990). Most existing works (Katz et al., 2017, 2019; Huang et al., 2017a, 2020; Wang et al., 2021b) utilize $L_{\infty}$ ball specifications for verifying robustness guarantees, i.e., verify if neural networks consistently predict the same label within local neighborhoods ($L_{\infty}$ balls) of reference data points.
Figure 1. The comparison between the NAP specification (depicted by the green region) and the $L_{\infty}$ ball specification (represented by the gray region) in the MNIST dataset. Image $a$ is the reference image, and image $c$ is the closest image to $a$ with an $L_{\infty}$ distance of 0.5294. However, image $c$ cannot be verified using the $L_{\infty}$ ball specification because an adversarial example $b$ can be found at an $L_{\infty}$ distance of 0.2. In contrast, the NAP specification enables verification of inputs that are semantically close despite their geometric distance being far apart, such as image $c$.
While this type of specification provides a fair and effective testbed for evaluating neural networks’ robustness, those verifiable $L_{\infty}$ balls are usually too tight, making them unlikely to cover unseen test set data (Geng et al., 2023a). For instance, the $L_{\infty}$ maximum verifiable bounds in the MNIST dataset is smaller than 0.2, whereas the smallest distance among images sharing the same label is greater than 0.5. Thus, narrowing the gap between verifiable regions and real data distribution to further enable verification in real-world settings poses a challenge.
To this end,  Geng et al. (2023a) proposes using neural activation patterns (NAPs) as specifications. Neural activation patterns can be viewed as an abstraction of each neuron’s state. For ReLU activation, it can be categorized into an activation state and a deactivation state. Since NAPs represent the intrinsic decision-making processes of neural networks, geologically distant data usually share the same NAPs if they belong to the same class. Compared to conventional local neighborhood specifications, NAP specifications cover much more flexible and potentially larger regions, as illustrated in Figure 1. This allows the verification of unseen real data rather than synthesized data from local neighbourhoods, paving the way for real-world application verification.
@@
However, it is noteworthy that the current approach to computing NAPs relies on a simple statistical method that assumes each neuron contributes to certifying the robustness behaviors of neural networks. Consequently, the computed NAPs are often the most refined ones. However, this is an overly restrictive assumption, as many studies (Frankle and Carbin, 2019; Lu et al., 2019; Liang et al., 2021; Wang et al., 2021a) have revealed that a non-trivial portion of neurons in neural networks may not play a substantial role. This motivates us to address the following challenge: Given a neural network, find a minimal (coarsest) NAP that is sufficient for formal verification of the network’s robustness. This problem is important for the following reasons: i) Minimal NAP specifications correspond to potentially much larger regions in the input space compared to the finest ones, this increases the potential of verification of unseen data; ii) Minimal NAP specifications provide insight into which neurons are responsible for the model making robust predictions, helping us to reveal the black-box nature of neural networks.
To find the minimal NAP specifications, we introduce two basic algorithms: Refine and Coarsen. These algorithms exhaustively check all possible candidates using off-the-shelf verification tools, such as Marabou (Katz et al., 2019). While these approaches provide correctness guarantees, they are not efficient for verifying large neural networks as calls to the verification tool are usually very expensive. To tackle this inefficiency, we propose statistical versions of Refine and CoarsenSample_Refine and Sample_Coarsen that leverage sampling and statistical learning principles to find the minimal NAP specification in a more efficient manner. However, current verification tools cannot scale up to state-of-the-art neural network models, limiting the ability of verification-dependent approaches to compute the minimal NAP specification.
This motivates us to study estimation methods that do not rely on verification tools. To this end, we find that adversarial examples and local gradients can provide valuable information about mandatory neurons, which are the essential building blocks of minimal NAPs. We then translate these findings into two approximate approaches: Adversarial_Prune and Gradient_Search. They can be used together to estimate mandatory neurons of state-of-art neural networks such as VGG-19 (Simonyan and Zisserman, 2014) and inspect possible links between models’ robustness performance and internal neurons, a task that existing verification frameworks struggle to scale for.
Moreover, previous research has claimed that NAP specifications cover significantly larger regions compared to $L_{\infty}$ ball specifications, yet this assertion lacks empirical evidence. In this paper, we address this gap by introducing a simple approach to estimate the volume of regions corresponding to NAP specifications. Furthermore, this method provides us with a rough approximation of the volumetric change from refined NAP specifications to minimal NAP specifications. Our experimental results demonstrate significant volumetric changes, thereby underscoring the efficacy of minimal NAP specifications in real-world verification scenarios. We summarize our contributions as follows:
  • We propose the problem of learning the minimal NAP specification and illustrate its importance in neural network verification. We present two simple approaches that leverage the verification tool with correctness guarantees.
  • We introduce several key concepts related to the problem of learning minimal NAP specifications, including the abstraction function, NAP specification, and mandatory neurons. We demonstrate that the problem can be easily solved with access to mandatory neurons and propose two efficient estimation approaches to learn them.
  • We propose two statistical learning approaches to the minimal NAP specification problem. They leverage sampling and estimation techniques as heuristics to achieve better efficiency.
  • We present a simple approach for estimating the volume of regions corresponding to NAPs. This helps us understand the volumetric change in verifiable space between the minimal NAP specifications and the refined NAP specifications.
  • Our experiments indicate that the minimal NAP specifications involve much smaller fractions of neurons compared to the most refined NAPs. And they significantly expand the verifiable bound by several orders of magnitude.
@@ 2.1. Neural Networks for Classification Tasks
In this paper, we focus on feed-forward ReLU neural networks. Generally speaking, a feed-forward network $N$ is comprised of $L$ layers, where each layer performs a linear transformation followed by a ReLU activation. We denote the pre-activation value and post-activation value at the $l$-th layer as $z^{(l)}(x)$ and $\hat{z}^{(l)}(x)$, respectively. The $l$-th layer computation is expressed as follows: $z^{(l)}(x)=\mathbf{W}^{(l)}\hat{z}^{(l-1)}(x)+\mathbf{b}^{(l)}$, $\hat{z}^{(l)}(x)=\mathbf{ReLu}(z^{(l)}(x))$, with $\mathbf{W}^{(l)}$ being the weight matrix and $\mathbf{b}^{(l)}$ representing the bias for the $l$-th layer. We denote the number of neurons in the $l$-th layer as $d_{l}$, and the $i$-th neuron in layer $l$ as $N_{i,l}$. The pre-activation value and post-activation value of $N_{i,l}$ at input $x$ is computed by $z_{i}^{(l)}(x)$ and $\hat{z}_{i}^{(l)}(x)$. The network can also be viewed as a function $\mathbf{F}^{<N>}:=\mathbb{R}^{d_{0}}\rightarrow\mathbb{R}^{d_{L}}$ such that $\mathbf{F}^{<N>}(x):=z^{(L)}(x)$, where $\mathbf{F}^{<N>}_{i}(x):=z_{i}^{(L)}(x)$ represents the output of $i$-th neuron in the last layer. We will omit $N$ when the context is clear. In a multi-class classification task with $C$ as the set of classes, when provided with an input $x$, the neural network predicts class/label $c\in C$ if the output of the $c$-th neuron $\textbf{F}_{c}(x)$ is the highest among the neurons in the $L$-th layer.
@@ 2.2. Adversarial Attacks and Neural Networks Verification Problem
Given a neural network $\mathbf{F}$ and a reference point $x$, adversarial attacks aim to search for a point $x^{\prime}$ that is geometrically close to the reference point $x$ such that $x^{\prime}$ and $x$ belong to different classes. Here, we use the canonical specification, that is, we want to search in the local neighborhoods ($L_{\infty}$ norm balls) of $x$, formally denoted as $B(x,\epsilon):=\{x^{\prime}\mid||x-x^{\prime}||_{\infty}\leq\epsilon\}$, where $\epsilon$ is the radius. For certain $\epsilon$, given we know $x$ belongs to class $j$, we say an adversarial point is found if:
(1) $\exists x^{\prime}\in B(x,\epsilon)\quad\exists i\in C\quad\mathbf{F}_{i}(x^{\prime})-\mathbf{F}_{j}(x)>0$
@@
Neural networks have been shown to be vulnerable to adversarial attacks. Even imperceptible changes could easily alter the predicted outcome. This highlights the importance of neural network verification. The neural network verification problem can be seen as the flip side of adversarial attacks. When it can be formally proven that no adversarial point exists in $B(x,\epsilon)$, the neural network verification problem is considered solved. Formally, we want to verify:
(2) $\forall x^{\prime}\in B(x,\epsilon)\quad\forall i\neq j\quad\mathbf{F}_{j}(x^{\prime})-\mathbf{F}_{i}(x^{\prime})>0$
For a simpler presentation, we assume that $\mathbf{F}(x)$ is a binary classifier. For any given specification, $\mathbf{F}(x)\geq 0$ indicates that the model is verified; otherwise, we can find an adversarial example. Solving such a problem is known to be NP-hard (Katz et al., 2017), and current state-of-the-art approaches rely on specialized solvers. The difficulty escalates significantly as the number of neurons in neural networks increases, limiting successful verifications to only very small neural networks. Therefore, achieving scalability in verification remains an ongoing challenge.
@@ 2.3. Neuron Abstractions and Neural Activation Patterns
To better discuss robustness verification using NAP specifications, we first introduce the relevant concepts of neuron abstractions, neuron abstraction functions, and neural activation patterns.
@@
[…]
@@
Given a neural network $N$ and the abstraction state set $\mathbb{S}$. A neuron abstraction function is the mapping $\mathcal{A}:N\rightarrow\mathbb{S}$. Formally, for an arbitrary neuron $N_{i,l}$ where $0\leq i\leq d_{l}$ and $11\leq l\leq L-$ 1, the function abstracts $N_{i,l}$ to some state $s_{k}\in\mathbb{S}$, i.e., $\mathcal{A}(N_{i,l})=s_{k}$.
The above characterization of neuron abstraction , it does not instruct us on how to perform binary abstraction in the absence of neuron values. Thus, we include the input(s) $x$ or $X$ as a new parameter in $\mathcal{A}$, and it is omitted when the context is clear. Here, we discuss three types of $\mathcal{A}$: the unary abstraction function function $\dot{\mathcal{A}}$, the binary abstraction function $\ddot{\mathcal{A}}$, and the statistical abstraction function $\widetilde{\mathcal{A}}$.
@@
The statistical $\widetilde{\mathcal{A}}$ function
The binary $\ddot{\mathcal{A}}$ function is limited in calculating a single input, and meets a challenge when multiple inputs come into play, as two distinct inputs may lead to disagreements in a neuron’s abstraction state. We then approach this problem statistically by introducing the $\widetilde{\mathcal{A}}$ function defined as follows:
where $\delta$ is a real number from $[0,1]$, and $X$ represents a set of inputs, i.e., $X:=\{x_{1},x_{2},...,x_{n}\}$. The parameter $\delta$ represents an error ratio to accommodate possible disagreements on neuron states between multiple inputs. Intuitively, the introduction of $\delta$ allows multiple inputs $x_{1},...,x_{n}$ to vote on a neuron’s state. For instance, when $\delta$ is set to 0.99, then 99% or more of the inputs must agree that a neuron is activated for the neuron to be in the $\mathbf{1}$ state.
(a) A simple 2x2 fully connected neuron network.
@@
Given a neural network $N$ and a neuron abstraction function $\mathcal{A}$. A neural activation pattern (NAP) $P$ is a tuple that consists of the abstraction state of all neurons in $N$. Formally, $P:=\langle\mathcal{A}(N_{i,l})\text{ }|\text{ }N_{i,l}\in N,\mathcal{A}\in\{\ddot{\mathcal{A}},\dot{\mathcal{A}}\}\rangle$, also denoted as $\mathcal{A}(N)$. The neuron $N_{i,l}$’s abstraction state specified by NAP $P$ is represented as $P_{i,l}$, i.e., $P_{i,l}=\mathcal{A}(N_{i,l})$.
Partially ordered NAP
@@
To give a concrete example, Figure 2(b) depicts a subset of NAPs of a simple neural network consisting of 2 hidden layers and 4 neurons, as presented in Figure 2(a).
@@
. This is because these NAPs sit on different branches formed by splitting certain ReLUs.
Region Outlined by NAPs
One key requirement for verification specification is its ability to represent a specific region within the input space.

### v4

# Learning Minimal Neural Specifications

@@ Abstract.
Formal verification is only as good as the specification of a system, which is also true for neural network verification. Existing specifications follow the paradigm of data as specification, where the local neighborhood around a reference data point is considered correct or robust. While these specifications provide a fair testbed for assessing model robustness, they are too restrictive for verifying any unseen test data points – a challenging task with significant real-world implications. Recent work shows great promise through a new paradigm, neural representation as specification, which uses neural activation patterns (NAPs) for this purpose. However, it computes the most refined NAPs, which include many redundant neurons. In this paper, we study the following problem: Given a neural network, find a minimal (general) NAP specification that is sufficient for formal verification of the network’s global robustness. Finding the minimal NAP specification not only expands verifiable bounds but also provides insights into which set of neurons contributes more to the model’s robustness. To address this problem, we propose three approaches—conservative, statistical, and optimistic—each offering distinct trade-offs between efficiency and performance. The first two rely on the verification tool to find minimal NAP specifications. The optimistic method efficiently estimates minimal NAPs using adversarial examples, without making calls to the verification tool until the very end. Each of these methods offers distinct strengths and trade-offs in terms of minimality and computational speed, making each approach suitable for scenarios with different priorities. The learnt minimal NAP specification allows us to inspect potential causal links between neurons and the robustness of state-of-the-art neural networks, a task for which existing work fails to scale. Our experimental results suggest that minimal NAP specifications require much smaller fractions of neurons compared to the NAP specifications computed by previous work, yet they can significantly expand the verifiable boundaries to several orders of magnitude larger.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
The growing prevalence of deep learning systems in decision-critical applications has elevated safety concerns regarding AI systems, such as their vulnerability to adversarial attacks  (Goodfellow et al., 2015; Dietterich and Horvitz, 2015). Therefore, the verification of AI systems has become increasingly important and attracted much attention from the research community. The field of neural network verification largely follows the paradigm of software verification – using formal methods to verify desirable properties of systems through rigorous mathematical specifications and proofs (Wing, 1990). A notable trend in existing works (Katz et al., 2017b, 2019; Huang et al., 2017a, 2020; Wang et al., 2021b) focuses on scaling verification to larger and more complex neural networks. While scalability is undeniably important and requires the collective effort of the research community, this paper explores an orthogonal angle that has been largely overlooked: defining meaningful specifications.
(a) Local neighborhood as specifications of the 3 reference points compared to using NAP as specification, NAP learnt from the 3 reference points and their local neighborhoods. Covers 4.23% and 7.64% of the WL region respectively.
(b) NAP as specification, NAP learnt from the 3 reference points. Covers 10.61% of the WL region.
(c) NAP as specification, NAP learnt from data points in the entire WL region. Covers 26.36% of the WL region.
(d) NAP as specification, with NAP learnt from the entire WL region and coarsened. Covers 32.43% of the WL region.
Figure 1. Illustration of verifiable regions of the Weak Left (WL) region of the ACAS_Xu (Katz et al., 2017a) advisories for a head-on encounter with $a_{\text{prev}}=\text{Clear of Conflict (COC)},\,\tau=0\text{s}$. The red bounding boxes represent the verifiable regions when using data as specifications, constrained by the reference points and their adversarial examples. The green hatched areas denote verifiable regions when using Neural Activation Patterns (NAP) as specifications.
To illustrate, nearly all existing works follow a ”data as specification” paradigm, where the specification is defined by the consistency of local neighborhoods—often $L_{\infty}$ balls—around reference data points. While local neighborhood specification provides a fair and effective testbed for evaluating neural network robustness, it primarily covers a tiny convex region of input data that can mathematically be described by adding noise to the reference point, as illustrated by the red bounding box in Figure 1(a). This small region is constrained by it’s adversarial examples. It is too restrictive to adequately address a broader, unseen test set, which are real data sampled from the underlying distribution.
Ideally, the specification should produce a verifiable region that includes all data points from the same class, where the data points may be distributed in a non-linear and non-convex manner within the input space. For instance, consider the safety-critical Airborne Collision Avoidance System for Unmanned Aircraft (ACAS Xu), the neural network processes an input representing the state of the aircraft and outputs one of five advisories: Clear of Conflict (COC), Weak Left (WL), Strong Left (SL), Weak Right (WR), or Strong Right (SR). Observe in 1 the the inputs belonging to Weak Left (WL) depicted in grey forms a irregular shape. This complexity underscores the need for flexible and robust specifications capable of accurately capturing such intricate distributions. Unlike local neighborhood specifications, manually defining such a specification is impractical. This poses a tricky chicken-egg problem: machine learning (ML) is necessary because it’s challenging to formally write down a precise definition (aka specification); but to be able to verify machine learning models, a formal specification would be needed. We argue that, in order to tackle this challenge, a separate learning algorithm for specifications is necessary. In this view, the “data as specification” paradigm is a simple but extremely overfitted algorithm for specification learning, which simply picks a small neighborhood of a reference data point in the input space. To this end, Geng et al. (2023) proposes using a new and more promising paradigm - ”neural representation as specification”, which learns a specification in the representation space of the trained machine learning model in the form of neural activation patterns (NAPs). NAPs are value abstractions of hidden neurons which have been shown useful for understanding the inference of a model (Gopinath et al., 2018). Most importantly, a well-trained neural network would exhibit similar activation patterns for input data from the same class, regardless of their actual distance in the input space (Bengio et al., 2013; Tishby and Zaslavsky, 2015; Geng et al., 2023). This key observation suggests that if we learn a NAP—a common activation pattern shared by a certain class of data—it can serve as a candidate specification for verifying data from that class. Once such a NAP specification is successfully verified, we say any data covered by this NAP (exhibiting this pattern) provably belongs to the corresponding class. Ideally, if that NAP covers all data from a class, it can be considered a machine-checkable definition of that class. Geometrically speaking, compared to the $L_{\infty}$ ball specifications, NAP specifications provide greater flexibility, enabling coverage of larger and more irregular regions, This advantage is particularly notable when reference points are near the boundaries of the WL region as illustrated in Figure 1(a). In 1(b) we observe that the common NAP learned from the three reference points covers a significantly larger region compared to the combined coverage of the individual NAPs learned from each reference point independently. 1(c) demonstrates even greater coverage when the NAP specification is learned from all inputs belonging to the WL class.
@@
However, it is noteworthy that the current approach to computing NAPs relies on a straightforward statistical method that assumes each neuron contributes to certifying the robustness of neural networks. Consequently, the computed NAPs are often overly refined. This is a restrictive assumption, as many studies (Frankle and Carbin, 2019; Lu et al., 2019; Liang et al., 2021; Wang et al., 2021a) have shown that a significant portion of neurons in neural networks may not play a substantial role. In the spirit of Occam’s Razor, we aim to systematically remove these redundant neurons that do not impact robustness. This motivates us to address the following challenge: given a neural network, how can we identify a minimal NAP (i.e., the coarsest abstraction) that is sufficient for verifying the network’s robustness? This problem is important for the following reasons: i) Minimal NAP specifications cover potentially much larger regions in the input space compared to the most refined ones, enhancing the ability to verify more unseen data; ii) Minimal NAP specifications provide insight into which neurons are responsible for the model’s robust predictions, helping to uncover the black-box nature of neural networks. For instance, if we aim to decode NAPs into human-understandable programs or rules (Li et al., 2023), minimal NAPs will always be easier to interpret than larger NAPs. We leave the interpretation of individual neurons as future work.
To find the minimal NAP specifications, we first introduce a basic algorithm, Coarsen, which exhaustively checks all possible candidates using off-the-shelf verification tools, such as Marabou (Katz et al., 2019). Specifically, Coarsen gradually coarsens each neuron in the most refined NAPs, retaining only the coarsened neurons when Marabou confirms verification success. While this approach provides correctness guarantees, it is not efficient for verifying large neural networks, as calls to verification tools are typically expensive. To improve efficiency, we further propose statistical variants of Coarsen—namely, StochCoarsen, which leverages sampling and statistical learning principles to find minimal NAP specifications. 1(d) highlights that coarsened NAP specifications, which involve fewer neurons, can cover an even larger portion of the WL region, highlighting the effectiveness of coarsened specifications in achieving broader coverage of the complex input space.
However, verification-dependent approaches face challenges in scaling up to state-of-the-art neural network models due to the limitations of current verification tools. To address this, we develop an optimistic method, OptAdvPrune, which leverages adversarial examples to identify essential neurons—the fundamental building blocks of minimal NAPs. Our experimental results show it can efficiently estimate an initial starting point for minimal NAP specification and only making calls to the verification tool until the very end. We further apply these methods to state-of-the-art neural networks, such as VGG-19 (Simonyan and Zisserman, 2014). While formal verification of these estimated minimal NAP specifications remains impractical due to current scalability constraints in the underlying verification engine, we demonstrate that these NAPs capture significant hidden features and concepts learned by the model. As many studies suggest, visual interpretability and robustness are inherently related, emerging in learned features and representations (Alvarez Melis and Jaakkola, 2018; Boopathy et al., 2020). Consequently, we believe that the identified essential neurons contribute to the model’s robustness, and their activation states can serve as empirical indicators of confident predictions. Our contributions can be summarized as follows:
  • We introduce the problem of learning minimal NAP specifications, emphasizing the need for a new paradigm in neural network specification. We present three approaches, each offering distinct trade-offs between efficiency and performance.
  • We define key concepts such as the abstraction function, NAP specification, and essential neurons. We demonstrate that the problem reduces to identifying essential neurons and propose an optimistic approaches for estimating them.
  • We propose a simple yet effective method for estimating the volume of the region mapped by the NAP. Our experiments show that minimal NAP specifications extend the verifiable bound by several orders of magnitude.
  • We estimate essential neurons in the state-of-the-art VGG-19 network. Using a modified Grad-CAM map, we show that these essential neurons contribute to visual interpretability, providing strong evidence that they may also explain the model’s robustness.
@@ 2.1. Neural Networks for Classification Tasks
In this paper, we focus on feed-forward neural networks equipped with ReLU activation functions. A feed-forward network $N$ consists of $L$ layers, where each layer alternates between a linear transformation and a ReLU activation. For the $l$-th layer, we denote the pre-activation and post-activation values as $z^{(l)}(x)$ and $\hat{z}^{(l)}(x)$, respectively. These are computed as follows: $z^{(l)}(x)=W^{(l)}\hat{z}^{(l-1)}(x)+b^{(l)},\quad\hat{z}^{(l)}(x)=ReLU(z^{(l)}(x))$, where $W^{(l)}$ and $b^{(l)}$ represent the weight matrix and bias vector of the $l$-th layer, respectively.
Let $d_{l}$ denote the number of neurons in the $l$-th layer, and let $N_{i,l}$ refer to the $i$-th neuron in that layer. The pre-activation and post-activation values of $N_{i,l}$ for a given input $x$ are $z_{i}^{(l)}(x)$ and $\hat{z}_{i}^{(l)}(x)$, respectively. The network as a whole can be viewed as a function $F^{<N>}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{L}}$, where $F^{<N>}(x)=z^{(L)}(x)$ represents the output of the final layer. For the $i$-th neuron in the last layer, its output is $F^{<N>}_{i}(x)=z_{i}^{(L)}(x)$. When the context is clear, we omit $N$ for simplicity and refer to the network as $F$. In multi-class classification tasks with $C$ classes, the network predicts the class $c\in C$ for a given input $x$ if the output $F_{c}(x)$ of the $c$-th neuron in the $L$-th layer is the largest among all outputs.
@@ 2.2. Adversarial Attacks and Neural Networks Verification Problem
Given a neural network $F$ and a reference point $x$, adversarial attacks aim to search for a point $x^{\prime}$ that is geometrically close to the reference point $x$ such that $x^{\prime}$ and $x$ belong to different classes. Here, we use the canonical specification, that is, we want to search in the local neighborhoods ($L_{\infty}$ norm balls) of $x$, formally denoted as $B(x,\epsilon):=\{x^{\prime}\mid||x-x^{\prime}||_{\infty}\leq\epsilon\}$, where $\epsilon$ is the radius. For certain $\epsilon$, given we know $x$ belongs to class $j$, we say an adversarial point is found if:
(1) $\exists x^{\prime}\in B(x,\epsilon)\quad\exists i\in C\quad F_{i}(x^{\prime})-F_{j}(x^{\prime})>0$
@@
Neural networks are vulnerable to adversarial attacks, where even imperceptible changes can alter predictions significantly. This underscores the critical need for neural network verification, which can be viewed as highlighting safe regions that excludes adversarial regions. Solving the verification problem involves formally proving the absence of adversarial points in $B(x,\epsilon)$. Formally, the verification problem seeks to prove:
(2) $\forall x^{\prime}\in B(x,\epsilon)\quad\forall i\neq j\quad F_{j}(x^{\prime})-F_{i}(x^{\prime})>0$
For a simpler presentation, we assume that $F(x)$ is a binary classifier. For any given specification $B(x,\epsilon)$, $F(x)\geq 0$ indicates that the model is verified; otherwise, we can find an adversarial example. Solving such a verification problem is known to be NP-hard (Katz et al., 2017b), and achieving scalability in verification remains an ongoing challenge.
@@ 2.3. Neuron Abstractions and Neural Activation Patterns
To better discuss NAP specification and robustness verification, we first introduce the relevant concepts of neuron abstractions, neuron abstraction functions, and neural activation patterns.
@@
For a neural network $N$, an internal neuron $N_{i,l}$ (where $0\leq i\leq d_{l}$ and $11\leq l\leq L-$ 1) can have its post-activation value $\hat{z}_{i}^{(l)}(x)$ abstracted into finite states. This abstraction is a mapping from $\mathbb{R}$ to a set $\mathbb{S}=\{s_{1},s_{2},\ldots,s_{n}\}$, representing neuron states. A simple binary abstraction for ReLU activation defines two states: $s_{0}:=0$ (deactivation) and $s_{1}:=(0,\infty)$ (activation). Further, these states can be abstracted into a unary state $s_{*}=[0,\infty)$, covering the entire range of post-activation. This leads to a partial order: $s_{*}\preceq s_{0}$ and $s_{*}\preceq s_{1}$, where $s_{0}$ and $s_{1}$ refine $s_{*}$. For simplicity, we use $\mathbf{0},\mathbf{1},\mathbf{*}$ to represent these states. Our specification learning approach applies to any activation function definable within the abstraction domain, though this study focuses on ReLU.
@@
Given a neural network $N$ and the abstraction state set $\mathbb{S}$. A neuron abstraction function is a mapping $\mathcal{A}:N\rightarrow\mathbb{S}$. Formally, for an arbitrary neuron $N_{i,l}$, the function abstracts $N_{i,l}$ to some state $s_{k}\in\mathbb{S}$, i.e., $\mathcal{A}(N_{i,l})=s_{k}$.
The above characterization of neuron abstraction does not instruct us on how to perform binary abstraction in the absence of neuron values. Therefore, we introduce two types of $\mathcal{A}$: unary $\dot{\mathcal{A}}$ and binary $\ddot{\mathcal{A}}$, both of which take a single input $x$ as a parameter, omitting it when the context is clear.
@@
(a) A simple 2x2 fully connected neural network.
@@
Given a neural network $N$, an input $x$, and neuron abstraction functions $\ddot{\mathcal{A}}$ and $\dot{\mathcal{A}}$, a neural activation pattern (NAP) $P$ is a tuple representing the abstraction state of all neurons in $N$. Formally, $P:=\langle\mathcal{A}(N_{i,l},x)\mid N_{i,l}\in N\rangle$, where $\mathcal{A}$ is either $\ddot{\mathcal{A}}$ or $\dot{\mathcal{A}}$, chosen for each neuron. We also denote $P$ as $\mathcal{A}(N,x)$. The abstraction state of neuron $N_{i,l}$ is represented as $P_{i,l}$, specifically $P_{i,l}=\mathcal{A}(N_{i,l},x)$. When representing a refined activation pattern triggered by $x$ where $\ddot{\mathcal{A}}$ is applied to each neuron, it is expressed as $\ddot{\mathcal{A}}(N,x)$.
@@
[…]
$\langle\mathbf{*},\mathbf{*},\mathbf{*},\mathbf{*}\rangle\preccurlyeq\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{*}\rangle\preccurlyeq\langle\mathbf{1},\mathbf{0},\mathbf{1},\mathbf{*}\rangle\preccurlyeq\langle\mathbf{1},\mathbf{0},\mathbf{1},\mathbf{0}\rangle.$
@@
. This occurs because these NAPs reside on different branches formed by splitting certain ReLUs.
Regions Outlined by NAPs